OpenStudy (anonymous):

f(x)=2x^2 PLEASE HELP

OpenStudy (anonymous):

what about it?

OpenStudy (anonymous):

I have no idea to do it. If it is the ABC method with the formula -b over 2a then im fine but i have never done a problem with just the A

OpenStudy (anonymous):

I think this is the equation.. (fx)=2x-2 y=2x-2 2x=y+2 x=(y+2)/2 f-¹(x)=(x+2)/2

OpenStudy (anonymous):

what do you mean you have no idea how to do it? You cannot "do" a function. Give the whole question.

OpenStudy (anonymous):

Complete parts a-c for each quadratic function. A. Find the y- intercept, the equation of the axis of symmetry, and the x-coordinate of the vertex.' B. Make a table of values that include the vertex. c. Use this info to graph the function.

OpenStudy (anonymous):

The y-intercept is 0. The function is symmetric around y-axis, so the equation is x=0. The x-coordinate of the vertex is 0.

OpenStudy (anonymous):



OpenStudy (anonymous):

Appreciate it!